mathslover (mathslover):

A dining hall has dimensions : 50 m * 15 m * 3.5 m . Calculate the mass of air in the hall ( density of air = 1.30 kg/m^3)

mathslover (mathslover):

@ghazi

OpenStudy (ghazi):

mass = volume *density = 50*15*3.5*1.30 kg this should be the answer but it doesnt seem to me reasonable :(

mathslover (mathslover):

@ghazi got it : \[\large{Density _ \textbf{air} = \frac{\textbf{mass of air}}{\textbf{Volume of air}}}\] \[\large{Density_\textbf{air} = \frac{\textbf{mass of air}}{\textbf{Volume of dining hall}}}\] so I got it now

mathslover (mathslover):

thanks @ghazi

mathslover (mathslover):

It seems reasonable to me :P

OpenStudy (ghazi):

you are welcome but always focus on the thing that your answer is reasonable

mathslover (mathslover):

:)

OpenStudy (ghazi):

:D